# Notions of entropy for unitary representations

### v2

@@ 9.2 Definition and first properties of almost periodic entropy
Let $\mathfrak{A}$ be a separable C*-algebra. Fix an AP sequence $\boldsymbol{\pi}=(\pi_{n})_{n\geq 1}$, a positive integer $k$, and an element $\varphi$ of $\mathfrak{B}(\mathfrak{A},\mathrm{M}_{k})_{+}$. Let $d_{n}$ be the dimension of $\pi_{n}$ for each $n$. Recall Definition 8.1 of the sets $\mathcal{X}(\pi_{n},O)$.
@@
$\mathrm{h}_{\boldsymbol{\pi}}(\varphi):=\inf_{O\ni\varphi}\limsup_{n\to\infty}\frac{1}{d_{n}}\log\frac{\mathrm{vol}_{2kd_{n}}\mathcal{X}(\pi_{n},O)}{v(d_{n})^{k}},$ (9.6)
@@
The division by $d_{n}$ outside the logarithm in (9.6) is a normalization choice, and is somewhat arbitrary. One could use $2d_{n}$ instead, but I have preferred to emphasize complex rather than real dimensions. Another natural choice would be $kd_{n}$, the complex dimension of the ambient space containing $\mathcal{X}(\pi_{n},O)$. However, the present choice seems to make for fewer explicit factors of $k$ later, for example in Lemma 9.7.
@@
Lemma 9.7 is analogous to the inequality between the entropy of a partition and the entropy rate of the generated process under a measure-preserving transformation: see, for instance, [Wal82, Theorem 4.12(i)]. We determine the cases of equality in Proposition 9.15 below.
For a given AP sequence $\boldsymbol{\pi}=(\pi_{n})_{n\geq 1}$ and completely positive map $\varphi$, the property of $\varphi$ being asymptotically associated to $\boldsymbol{\pi}$ depends only on the approximate equivalence class of $\pi_{\varphi}$. However, the actual value $\mathrm{h}_{\boldsymbol{\pi}}(\varphi)$ is more sensitive. This is an important point where our story diverges from ergodic theory. In ergodic theory, one of the most essential properties of sofic entropy is its independence of the choice of generating observable [Bow10b], and hence its invariance under isomorphism for abstract measure-preserving systems. The analog of this for AP entropy is false. Even if we restrict attention to single cyclic unit vectors in a representation $\pi$, the resulting values of AP entropy can vary greatly.
Rather than invariance, AP entropy enjoys a general transformation law when one cyclic tuple is exchanged for another: see Proposition 9.10.
In some cases, proofs about AP entropy are much easier to digest in the special case $k=1$, if only because the notation is lighter. The next result sometimes lets us make this simplification without losing any generality. It accompanies Lemmas 8.15 and 8.18.
@@
Let $k$ be a positive integer, let $\boldsymbol{\pi}=(\pi_{n})_{n\geq 1}$ be an AP sequence, and let $\boldsymbol{\pi}^{(k)}:=(\pi_{n}^{(k)})_{n\geq 1}$. Let $\varphi\in\mathfrak{B}(\mathfrak{A},\mathrm{M}_{k})_{+}$ be completely positive, and define $\langle\varphi,\cdot\rangle$ as in equation (4.3). Then
@@
The Introduction discusses AP entropy as a representation theoretic analog of sofic entropy. Let us comment further on their relationship here.
Theorem C applies to an AP sequence that is assumed to trace-converge to some $\tau$. As a result, the Fuglede–Kadison determinant that appears in Theorem C can arise as an AP entropy value whenever $\tau$ is a limit of tracial states pulled back from some AP sequence.
Suppose that $\mathfrak{A}=C^{\ast}\Gamma$ and that $\tau$ comes from the regular charater. Then it is a limit of finite-dimensional characters if and only if $\Gamma$ has an AP sequence $(\pi_{i})_{i\geq 1}$ that separates its elements. Such a group is called ‘maximally almost periodic’. For finitely presented groups this property turns out to be equivalent to residual finiteness. Maximally almost periodic groups were introduced by von Neumann [vN34] and covered by Weil in [Wei40, Chapter VII] and Dixmier in [Dix77, Sections 16.4–5] (where they are called ‘injectable’ groups).
However, we can also study AP entropy for regular characters on a larger class of groups as follows. If $\Gamma$ is any countable group, then we can write it as $F/N$ for some free group $F$ and normal subgroup $N$. Now we can look for finite-dimensional representations of $F$ whose characters converge to the quasi-regular character $1_{N}$, rather than finite-dimensional representations of $\Gamma$ whose characters converge to $1_{\{e\}}$. This offers more flexibility, because those finite-dimensional representations of $F$ need not have trivial restriction to $N$ until we take their limit. Allowing convergence in this sense, the availability of finite-dimensional approximants to the regular representation of $\Gamma$ is equivalent to $\Gamma$ being ‘hyperlinear’ in the terminology of [R0̆8]. This class of groups is the ‘linear’ analog of the sofic groups, and includes the sofic groups by another result from [ES05]. Analogously, soficity itself can be formulated in terms of quotients of a free group and finitary approximations to the resulting Schreier graphs; see [Can15, Section 4], for example.
@@
See [BdlH20, Subsection 4.C.b] for a more modern introduction to maximally almost periodic groups, and [BO08, Section 6.4.2] for a connection to Kirchberg’s ‘factorization property’.
@@ 9.3 First transformation formula
Fix an AP sequence $\boldsymbol{\pi}=(\pi_{n})_{n\geq 1}$, and let $d_{n}$ be the dimension of $\pi_{n}$. At this point, our work starts to need the assumption that the tracial states $\overline{\mathrm{tr}}_{d_{n}}\circ\pi_{n}$ converge to a limit in $\mathfrak{A}^{\ast}$.
@@
    $\psi(b):=Q^{\ast}\varphi(b)Q\qquad(b\in\mathfrak{A}).$
@@
    Assume further that $\overline{\mathrm{tr}}_{d_{n}}\circ\pi_{n}\to\tau$. Let $a\in\mathrm{M}_{k}(\mathfrak{A})$ be invertible, and define $\psi\in\mathfrak{B}(\mathfrak{A};\mathrm{M}_{k})_{+}$ in terms of $\varphi$ and $a$ as in equation (4.6) of Lemma 4.5. Then $\varphi$ is asymptotically associated to $\boldsymbol{\pi}$ if and only if $\psi$ is, and in that case
@@
Part (a). For any neighbourhood $O$ of $\psi$, Lemma 8.7 gives a neighbourhood $U$ of $\varphi$ such that
$\mathcal{X}(\pi,U)\cdot Q\subset\mathcal{X}(\pi,O)$
for any representation $\pi$. For each $n$, it follows that
$\mathrm{vol}_{2kd_{n}}\mathcal{X}(\pi_{n},O)\geq|\det Q|^{2d_{n}}\cdot\mathrm{vol}_{2kd_{n}}\mathcal{X}(\pi_{n},U).$
The determinant is raised to the power $2d_{n}$ here because $Q$ is a $k$-by-$k$ complex linear transformation, but we are using it to define a linear transformation in $kd_{n}$ complex dimensions, and we must treat it as a real linear map in $2kd_{n}$ real dimensions for the purpose of computing volumes (see, for instance, [Rud08, Subsection 1.3.5]). Inserting this inequality into Definition 9.4, it follows that
@@
Part (b). If $O$ is a neighbourhood of $\psi$, then Lemma 8.5 gives a neighbourhood $O^{\prime}$ of $\varphi$ such that
@@
The reverse of this inequality also holds by swapping the roles of $\varphi$ and $\psi$ and replaced $a$ with $a^{-1}$, which satisfies
@@ 9.4 Spherical measures and concentration
for some $O\in\mathcal{O}$ and $\delta>0$. Then $\mathcal{U}$ is a base of neighbourhoods around $\varphi$ in $\mathfrak{A}^{\ast}$. We may therefore restrict attention to neighbourhoods from $\mathcal{U}$ when evaluating $\mathrm{h}_{\boldsymbol{\pi}}(\varphi)$ (see Remark 9.5).
@@
As a result, integrating in polar coordinates [Rud08, Subsection 1.4.3] gives
@@
for all sufficiently large $d$. Inserting this into Definition 9.4 and taking the limit supremum over $n$ and then the infimum over $O$ and $\delta$, we obtain (9.8). ∎
If $\varphi\in\Sigma_{k}(\mathfrak{A})$ for some $k>1$, then we may apply Lemma 9.12 to the positive functional $\langle\varphi,\cdot\rangle$ on $\mathrm{M}_{k}(\mathfrak{A})$, whose sets of approximately typical vectors along $\boldsymbol{\pi}^{(k)}$ are given by (8.2). We can then deduce a result for $\varphi$ itself via Lemma 9.8.
For a state, we now have a choice between the original definition of AP entropy and the alternative given by Lemma 9.12. Each has its advantages. The first major advantage of the measures $\sigma_{2d-1}$ is contact with the measure concentration from Lemma 9.1. As above, we discuss this only for $k=1$ to lighten notation. Let $\pi$ be a $d$-dimensional representation. Since the trace of a matrix is invariant under unitary conjugation, the average of the type $\Phi^{\pi}_{x}$ with respect to the spherical measure $\sigma_{2d-1}$ is equal to $\overline{\mathrm{tr}}_{d}\circ\pi$. When $d$ is large, Lemma 9.1 improves this conclusion considerably: $\Phi^{\pi}_{x}$ is actually close to $\overline{\mathrm{tr}}_{d}\circ\pi$ for most individual $x\in\mathrm{S}^{2d-1}$.
@@
Since $\overline{\mathrm{tr}}_{d_{n}}\circ\pi_{n}\to\tau$, this $U$ satisfies
@@
We may assume that $\varphi$ is normalized. Proposition 4.7 gives that $\Sigma_{k}(\lambda^{\oplus\infty})$ equals $\Sigma_{k}(\lambda^{\oplus k})$, so now our assumption actually says that $\varphi\in\overline{\Sigma_{k}(\lambda^{\oplus k})}$. Therefore, by Corollary 8.6, it suffices to show that $\tau\otimes I_{k}$ itself is asymptotically associated to $\boldsymbol{\pi}$ (recall Example 4.1 for this tensor-product notation). When $k=1$, this follows from Lemma 9.13: indeed, a random vector drawn from $\sigma_{2d_{n}-1}$ is approximately typical for $\tau$ with high probability once $d_{n}$ is large enough. Finally, if $k>1$, then we can apply the previous case to the sequence $\boldsymbol{\pi}^{(k)}$ using Lemmas 8.18 and 8.15 and the fact that
@@
In the notation of Subsection 8.3, Corollary 9.14 gives a lower bound on the set $\mathrm{T}\liminf_{n}\overline{\Sigma_{k}(\pi_{n})}$ for each $k$. This can be an equality: for example, this is so if $\mathfrak{A}=C^{\ast}\Gamma$, the states $\overline{\mathrm{tr}}_{d_{n}}\circ\pi_{n}$ converge to the regular tracial state, and $(\pi_{n})_{n\geq 1}$ also converges strongly to the regular representation. Some important examples satisfying these three conditions are surveyed in [Mag25]. But in general the inclusion can be strict.
We can now use Lemmas 9.12 and Lemma 9.13 to determine the cases of equality in Lemma 9.7.
@@
On the other hand, suppose that $\varphi\neq\varphi(1)\cdot\tau$. After normalizing using Proposition 9.10(a), we may assume that $\varphi(1)=1$, and so Lemma 9.12 applies to $\varphi$.
Since $\varphi\neq\tau$, they have disjoint neighbourhoods, say $U$ and $U^{\prime}$ respectively. Applying Lemma 9.13 to the neighbourhood $U^{\prime}$, there are positive constants $C$ and $c$ such that
@@
The definition of $\varphi$ is a special case of equation (4.6) relating the maps $\tau$ and $\varphi$ via $a$. Therefore Proposition 9.10 gives

### v3

@@ 9.2 Definition and first properties of almost periodic entropy
Consider again a separable, unital C*-algebra $\mathfrak{A}$. Fix an AP sequence $\boldsymbol{\pi}=(\pi_{n})_{n\geq 1}$, a positive integer $k$, and an element $\varphi$ of $\mathfrak{B}(\mathfrak{A},\mathrm{M}_{k})_{+}$. Let $d_{n}$ be the dimension of $\pi_{n}$ for each $n$. Recall Definition 8.1 of the sets $\mathcal{X}(\pi_{n},O)$.
@@
$\mathrm{h}_{\boldsymbol{\pi}}(\varphi):=\inf_{O}\limsup_{n\to\infty}\frac{1}{d_{n}}\log\frac{\mathrm{vol}_{2kd_{n}}\mathcal{X}(\pi_{n},O)}{v(d_{n})^{k}},$ (9.6)
@@
The constant outside the logarithm in (9.6) is a normalization choice, and is somewhat arbitrary. One could divide by $2d_{n}$ instead, but I have preferred to emphasize complex rather than real dimensions. Another natural choice would be $kd_{n}$, the complex dimension of the ambient space containing $\mathcal{X}(\pi_{n},O)$. However, the present choice seems to make for fewer explicit factors of $k$ later, for example in Lemma 9.7 below.
@@
Lemma 9.7 is analogous to the inequality between the entropy of a partition and the entropy rate of the generated process under a measure-preserving transformation: see, for instance, [157, Theorem 4.12(i)]. We determine the cases of equality in Proposition 9.17 below.
For a given AP sequence $\boldsymbol{\pi}=(\pi_{n})_{n\geq 1}$ and completely positive map $\varphi$, the property of $\varphi$ being asymptotically associated to $\boldsymbol{\pi}$ depends only on the equivalence class of $\pi_{\varphi}$. However, the actual value $\mathrm{h}_{\boldsymbol{\pi}}(\varphi)$ is more sensitive. This is an important point where our story diverges from ergodic theory. In ergodic theory, one of the most essential properties of sofic entropy is its independence of the choice of generating observable [25], and hence its invariance under isomorphism for abstract measure-preserving systems. The analog of this for AP entropy is false. Even if we restrict attention to single cyclic unit vectors in a representation $\pi$, the resulting values of AP entropy can vary greatly.
However, rather than invariance, AP entropy enjoys a general transformation law when one cyclic tuple is exchanged for another: see Proposition 9.10.
Some proofs about AP entropy are easier to digest in the special case $k=1$, if only because the notation is lighter. The next result sometimes lets us make this simplification without losing any generality. It accompanies Lemmas 8.11 and 8.15.
@@
Let $k$ be a positive integer, let $\boldsymbol{\pi}=(\pi_{n})_{n\geq 1}$ be an AP sequence, and let $\boldsymbol{\pi}^{(k)}:=(\pi_{n}^{(k)})_{n\geq 1}$. Let $\varphi\in\mathfrak{B}(\mathfrak{A},\mathrm{M}_{k})_{+}$, and define $\langle\varphi,\cdot\rangle$ as in equation (4.7). Then
@@
The Introduction discusses AP entropy as a representation theoretic analog of sofic entropy. Let us comment further on their relationship.
Theorem C applies to an AP sequence whose pulled-back traces converge to some $\tau$. As a result, the Fuglede–Kadison determinant that appears in Theorem C can arise as an AP entropy value whenever $\tau$ is a limit of tracial states pulled back from some AP sequence.
Suppose that $\mathfrak{A}=C^{\ast}\Gamma$ and that $\tau$ comes from the regular charater. Then it is a limit of finite-dimensional characters if and only if $\Gamma$ has an AP sequence $(\pi_{i})_{i\geq 1}$ that separates its elements. Such a group is called ‘maximally almost periodic’. For finitely presented groups this property turns out to be equivalent to residual finiteness. Maximally almost periodic groups were introduced by von Neumann [150] and covered by Weil in [158, Chapter VII] and Dixmier in [48, Sections 16.4–5] (where they are called ‘injectable’ groups).
However, we can also study AP entropy for regular characters on a larger class of groups as follows. If $\Gamma$ is any countable group, then we can write it as $F/N$ for some free group $F$ and normal subgroup $N$. Now we can look for finite-dimensional representations of $F$ whose characters converge to the quasi-regular character $1_{N}$, rather than finite-dimensional representations of $\Gamma$ whose characters converge to $1_{\{e\}}$. This offers more flexibility, because those finite-dimensional representations of $F$ need not have trivial restriction to $N$ until we take their limit. Allowing convergence in this sense, the availability of finite-dimensional approximants to the regular representation of $\Gamma$ is equivalent to $\Gamma$ being ‘hyperlinear’ in the terminology of [125]; in particular, it does not depend on the choice of presentation $F/N$. This class of groups is the ‘linear’ analog of the sofic groups, and includes the sofic groups by a result from [53]. Analogously, soficity itself can be formulated in terms of quotients of a free group and finitary approximations to the resulting Schreier graphs; see [32, Section 4], for example.
@@
See [14, Subsection 4.C.b] for a modern introduction to maximally almost periodic groups, and [30, Section 6.4.2] for a connection to Kirchberg’s ‘factorization property’.
@@ 9.3 First transformation formula
Fix an AP sequence $\boldsymbol{\pi}=(\pi_{n})_{n\geq 1}$, and let $d_{n}$ be the dimension of $\pi_{n}$. At this point, some of our work starts to need the assumption that the tracial states $\overline{\mathrm{tr}}_{d_{n}}\circ\pi_{n}$ converge to a limit in $\mathfrak{A}^{\ast}$.
The next proposition is a basic change-of-variables formula for AP entropy. It can be seen as a cousin of Voiculescu’s change-of-variables formulas for free entropy in free probability theory [154, 155], although the proof in our ‘linear’ setting is simpler.
@@
    $\psi(b):=(Q^{\mathrm{T}})^{\ast}\varphi(b)Q^{\mathrm{T}}\qquad(b\in\mathfrak{A}).$
@@
    Assume further that $\overline{\mathrm{tr}}_{d_{n}}\circ\pi_{n}\to\tau$. Let $a\in\mathrm{M}_{k}(\mathfrak{A})$ be invertible, and define $\psi\in\mathfrak{B}(\mathfrak{A};\mathrm{M}_{k})_{+}$ in terms of $\varphi$ and $a$ as in Lemma 4.3. Then $\varphi$ is asymptotically associated to $\boldsymbol{\pi}$ if and only if $\psi$ is, and in that case
@@
Part (a). For any neighbourhood $O$ of $\psi$, the second part of Lemma 8.4 gives a neighbourhood $U$ of $\varphi$ such that
$(I_{d}\otimes Q)[\mathcal{X}(\pi,U)]\subset\mathcal{X}(\pi,O)$
for any $d$-dimensional representation $\pi$. For each $n$, it follows that
$\mathrm{vol}_{2kd_{n}}\mathcal{X}(\pi_{n},O)$ $\geq|\det(I_{d_{n}}\otimes Q)|^{2}\cdot\mathrm{vol}_{2kd_{n}}\mathcal{X}(\pi_{n},U)$
$=|\det Q|^{2d_{n}}\cdot\mathrm{vol}_{2kd_{n}}\mathcal{X}(\pi_{n},U).$
The determinants are squared here because $I_{d_{n}}\otimes Q$ is a linear transformation in $kd_{n}$ complex dimensions, but we must treat it as a real linear map in $2kd_{n}$ real dimensions for the purpose of computing volumes (see, for instance, [127, Subsection 1.3.5]). Inserting this inequality into Definition 9.4 and taking the infimum over $O$, it follows that
@@
Part (b). This time, if $O$ is a neighbourhood of $\psi$, then Lemma 8.4 gives a neighbourhood $O^{\prime}$ of $\varphi$ such that
@@
The reverse of this inequality also holds by swapping the roles of $\varphi$ and $\psi$ and replacing $a$ with $a^{-1}$, which satisfies
@@ 9.4 Spherical measures and concentration
for some $O\in\mathcal{O}$ and $\delta>0$. Then $\mathcal{U}$ is a base of neighbourhoods around $\varphi$ in $\mathfrak{A}_{+}^{\ast}$. We may therefore restrict attention to neighbourhoods from $\mathcal{U}$ when evaluating $\mathrm{h}_{\boldsymbol{\pi}}(\varphi)$ (see Remark 9.5).
@@
As a result, integrating in polar coordinates [127, Subsection 1.4.3] gives
@@
for all sufficiently large $d$. Inserting this into Definition 9.4 and taking the limit supremum over $n$ and then the infimum over elements if $\mathcal{U}$, we obtain (9.8). ∎
Remark 9.13.
[…]
If $\varphi\in\Sigma_{k}(\mathfrak{A})$ for some $k>1$, then we may apply Lemma 9.12 to the positive functional $\langle\varphi,\cdot\rangle$ on $\mathrm{M}_{k}(\mathfrak{A})$, whose sets of approximately typical vectors along $\boldsymbol{\pi}^{(k)}$ are given by equation (8.2). We can then deduce a result for $\varphi$ itself via Lemma 9.8.
For a state, we now have a choice between the original definition of AP entropy and the alternative given by Lemma 9.12. Each has its advantages. A major advantage of the measures $\sigma_{2d-1}$ is contact with the measure concentration from Lemma 9.1. As above, we discuss this only for $k=1$ to lighten notation. Let $\pi$ be a $d$-dimensional representation. Since the trace of a matrix is invariant under unitary conjugation, the average of the type $\Phi^{\pi}_{x}$ with respect to the spherical measure $\sigma_{2d-1}$ is equal to $\overline{\mathrm{tr}}_{d}\circ\pi$. When $d$ is large, Lemma 9.1 improves this conclusion considerably: $\Phi^{\pi}_{x}$ is actually close to $\overline{\mathrm{tr}}_{d}\circ\pi$ for most individual $x\in\mathrm{S}^{2d-1}$.
@@
Since $\overline{\mathrm{tr}}_{d_{n}}\circ\pi_{n}(a)\to\tau(a)$, this $U$ satisfies
@@
We may assume that $\varphi$ is normalized. Proposition 4.10 gives that $\Sigma_{k}(\lambda^{\oplus\infty})$ equals $\Sigma_{k}(\lambda^{\oplus k})$, so now our assumption actually says that $\varphi\in\overline{\Sigma_{k}(\lambda^{\oplus k})}$.
Since $\lambda^{\oplus k}$ is the minimal dilation of $\tau\otimes I_{k}$, by Corollary 8.5 it suffices to show that $\tau\otimes I_{k}$ itself is asymptotically associated to $\boldsymbol{\pi}$ (recall Example 4.1 for this tensor-product notation). When $k=1$, this follows from Lemma 9.14: indeed, a random vector drawn from $\sigma_{2d_{n}-1}$ is approximately typical for $\tau$ with high probability once $d_{n}$ is large enough. Finally, if $k>1$, then we can apply the previous case to the sequence $\boldsymbol{\pi}^{(k)}$ using Lemmas 8.15 and 8.11 and the fact that
@@
In the notation of Subsection 8.3, Corollary 9.15 gives a lower bound on the set $\mathrm{T}\liminf_{n}\overline{\Sigma_{k}(\pi_{n})}$ for each $k$. This can be an equality: for example, this is so if $\mathfrak{A}=C^{\ast}\Gamma$, the states $\overline{\mathrm{tr}}_{d_{n}}\circ\pi_{n}$ converge to the regular tracial state, and $(\pi_{n})_{n\geq 1}$ also converges strongly to the regular representation. Some important examples satisfying these three conditions are surveyed in [113]. But in general the inclusion can be strict. Nevertheless, we do always obtain the following.
Corollary 9.16.
Assume that $\overline{\mathrm{tr}}_{d_{n}}\circ\pi_{n}\to\tau$. Let $\varphi\in\mathfrak{B}(\mathfrak{A},\mathrm{M}_{k})_{+}$, and consider its Lebesgue decomposition relative to $\tau$ as given by Proposition 4.9. Then $\varphi$ is asymptotically associated to $\boldsymbol{\pi}$ if and only if $\varphi_{\mathrm{sing}}$ is asymptotically associated to $\boldsymbol{\pi}$.
Proof.
First, by Proposition 4.9 and Lemmas 8.11 and 8.15, it suffices to prove that $\langle\varphi,\cdot\rangle$ is asymptotically associated to $\boldsymbol{\pi}^{(k)}$ if and only if $\langle\varphi_{\mathrm{sing}},\cdot\rangle$ is asymptotically associated to $\boldsymbol{\pi}^{(k)}$. After adjusting some notation if necessary, this reduces our work to the case when $k=1$. Assume this for the rest of the proof.
By Proposition 3.2, $\varphi_{\mathrm{sing}}$ is associated to $\pi_{\varphi}$. Therefore, if $\varphi$ is asymptotically associated to $\boldsymbol{\pi}$, then $\varphi_{\mathrm{sing}}$ is as well by Corollary 8.5.
On the other hand, Corollary 9.15 tells us that $\varphi_{\mathrm{ac}}$ is always asymptotically associated to $\boldsymbol{\pi}$, and $\varphi_{\mathrm{ac}}$ and $\varphi_{\mathrm{sing}}$ are disjoint by construction. Therefore, if $\varphi_{\mathrm{sing}}$ is asymptotically associated to $\boldsymbol{\pi}$, then so is $\varphi$ itself by Lemma 8.12. ∎
Lemmas 9.12 and 9.14 let us determine the cases of equality in Lemma 9.7.
@@
On the other hand, suppose that $\varphi\neq\varphi(1)\cdot\tau$. After normalizing using Proposition 9.10(a), we may assume that $\varphi(1)=1$. Since $\varphi\neq\tau$, they have disjoint neighbourhoods, say $U$ and $U^{\prime}$ respectively. Applying Lemma 9.14 to the neighbourhood $U^{\prime}$, there are positive constants $C$ and $c$ such that
@@
The expression for $\varphi$ in terms of $\tau$ and $a$ is a special case of the relationship from Lemma 4.3. Therefore Proposition 9.10(b) gives